OpenStudy (anonymous):

A focus of an ellipse is at the origin. The directrix is the line x = 4 and the eccentricity is 1/2. Then thelength of the semi−major axis is

OpenStudy (anonymous):

@experimentX @hartnn @atlas @Psymon @zepdrix

OpenStudy (psymon):

Lol, never seen directrix used for an ellipse, so Im all dumb xD

OpenStudy (anonymous):

@satellite73 @terenzreignz @.Sam. @experimentX

OpenStudy (anonymous):

@atlas

OpenStudy (experimentx):

let center by (x1, 0) and the ellipse cut x-axis at (x2,0). your length of semi major axis is (x2 -x1) you know that foci = ae = |(x2 - x1)| 1/2 = |x1| and 1/2 |(4-x2)| = |x2 - x1| solve this two equations and get |x2-x1| this is your semi major axis.